# Multiplexing in Networks and Diffusion

### v1

@@ 2.2.2. Correlations among Layers
Figure 1 reveals several patterns. First, there are consistently high correlations between layers in both data sets—above 0.5 for most layer pairs. Second, the exceptions are the distance, jati, and temple layers. The jati and distance layers are almost uncorrelated with the other layers, [^8] This does not mean, for instance, that there is not substantial jati-based homophily in these data. The low correlation comes from the fact that the jati layer dramatically over-predicts relationships compared to other layers, so it has many 1’s where there are 0’s in the other layers., [^9] Distance is higher when people live far from each other and are thus less likely to be linked, all else held equal; this explains the negative signs. while the temple layer has an intermediate level of correlation with others. Third, the layers are more highly correlated in the RCT villages compared to the microfinance villages.
2.2.3. Principal Component Analyses
Our third look at the structure of the networks uses principal component analyses, conducted in two stages.
First, we perform a principal component analysis with all of the layers (excluding the synthetic union and intersection layers). We treat each pair of households (in a given village) as an observation, yielding $\sum_{v}\binom{n_{v}}{2}$ observations, where $n_{v}$ is the number of households in village $v$, and the number of dimensions is the number of layers $L$ in the given sample.
@@ 2.2.3. Principal Component Analyses
As we see in Figure 2, when including all layers, the first principal component aligns with most relationship layers, capturing almost half (48.7%) of the variation in the microfinance villages and more than two-thirds (72.1%) in the RCT villages. Panels A and B plot the coordinates of the first and second component entries for each link type. Interestingly, jati largely aligns with the second principal component, as one would expect given its relatively low correlation with the other layers. The geographic distance layer is nearly opposite to jati, which reflects the fact that people from the same jati live in close proximity. The complete results appear in Appendix Tables S2 and S3.
Next, in Figure 3 we repeat the analysis after removing the least correlated dimensions: jati, geography, and temple.101010 We also redo the analysis just dropping jati and geography and keeping temple in Supplemental Appendix Figure S2. Temple is sparse and essentially orthogonal to the other dimensions. This allows us to zoom in on the correlation patterns among the social and economic layers.
@@
Building the Backbone.
To capture the correlation structure of the network layers, we use the principal component analysis to construct an aggregate network from the multigraph, which we call the backbone. The backbone network is built using the first $K$ principal components, constructed as described above. The $K$ we use is determined by a so-called ladle plot, with the goal of selecting a cutoff yielding an “optimal” low-dimensional representation (see Luo and Li (2016) for details). [^11] To select the optimal number of principal components the literature usually relied on a cutoff based on patterns of either decreasing eigenvalues or increasing variability of eigenvectors. Luo and Li (2016) combine these two approaches to better estimate the optimal K. They propose a new estimator, called the “ladle estimator” which minimizes an objective function that incorporates both the magnitude of eigenvalues and the bootstrap variability of eigenvectors. This approach exploits the pattern that when eigenvalues are close together, their corresponding eigenvectors tend to vary greatly, and when eigenvalues are far apart, the eigenvector variability tends to be small. By leveraging both sources of information, the ladle estimator can more precisely determine the rank of the matrix, and thus the optimal number of components to retain.
$w_{k}:=\lambda_{k}/\sum_{j=1}^{K}\lambda_{j}.$
For each village $v$, we then define a “backbone” network, $g^{text{backbone}}$, from the principal components as a weighted graph where
$g_{ij,v}^{\text{backbone}}=Z_{ij,v}.$
In words, the backbone reduces the multiplex data to a synthetic structure by projecting the multidimensional links onto the top principal components. After determining the number of components $K$, we compute each dyad’s coordinates along these components, scaling by eigenvalues, which (as is standard in PCA) quantify the importance of various dimensions. Summing these weighted projections yields a single index.
2.3. Determinants of Diffusion
@@
$DC_{j,v}^{\ell}:=\left[\sum_{t}^{T}(qg^{\ell}_{v})^{t}\cdot 1\right]_{j},$
$DC_{v}^{\ell}:=\sum_{j\in S_{v}}DC_{j,v}^{\ell}.$
@@
Table 2 depicts how differently the layers predict diffusion based on our specification in (2.1). (Supplementary Appendix Figure S1 plots the $90$% and $95$% confidence intervals.)
@@
Interestingly, the four synthetic networks we have mentioned that aggregate the layers in specific ways—union, intersection, total, and backbone—all perform worse than the individual layers with the exception of jati. However, this appears to be rooted in the inclusion of jati in those aggregates. In Supplement A.2 we recreate Table 2 with aggregate layers that omit jati in their construction (this applies only to union, intersection, and backbone). This improves their performance, with the backbone network now yielding an $R^{2}$ second only to the advice layer.
@@
(2.2) $y_{v}=\alpha+\sum_{\ell}\beta^{\ell}\cdot DC_{v}^{\ell}+X_{v}\Gamma^{\ell}+\epsilon_{v,\ell}$
@@
A complication we face here is that in order to be consistent, LASSO requires a condition called irrepresentability, which requires the regressors of interest not to be excessively correlated (Zhao and Yu, 2006). In our setting, this requirement fails since the network layers are highly correlated. To overcome this problem, we use the Puffer transformation developed by Rohe (2015) and Jia et al. (2015), which recovers irrepresentability when the number of observations exceeds the number of variables. Although the regressors $(DC^{\ell}_{v})_{v,\ell}$ have correlated columns, by appropriately pre-conditioning the data matrix, we can force its columns to be orthogonal and therefore irrepresentable. Puffer-LASSO then recovers the set of relevant variables with probability tending to one exponentially fast in the number of observations, with consistent parameter estimates, that are asymptotically normally distributed with probability approaching one (Javanmard and Montanari, 2013; Jia et al., 2015; Taylor and Tibshirani, 2015; Lee et al., 2016; Banerjee et al., 2024a).
@@
Table 4 presents both cumulative and marginal F-tests as variables are added in the order selected by LASSO. We can see that adding intersection is marginally significant above advice, and further including kerorice and jati yields a more complete model, with an improvement significant at the 5 percent level.131313 In Appendix Table S4 we exclude the extra layers of intersection, union, and backbone, which are “constructed” layers that are derived from these basic layers. F-tests include the basic layers in the order selected by the Lasso. Thus, even though jati serves as a poor substitute for other layers, it turns out to be a useful complement to them in predicting diffusion.
@@ 2.4. How the Level of Multiplexing Affects Diffusion
The multiplexing score for a household $i$ measures the average fraction of relationship types it has with each of its neighbors. The numerator calculates the average number of links household $i$ has to each neighbor across all $L$ relationship types. It does this by first summing the number of links between household $i$ and each neighbor $j$ across all layers, dividing by the total number of layers $L$, and then summing this average across all neighbors $j$. The denominator counts the number of unique neighbors of household $i$ by summing an indicator for whether there is at least one link between $i$ and $j$ across any layer. For example, $m_{i,v}=1$ if whenever household $i$ has a relationship with some other household $j$, then it has all possible relationships with that other household. In contrast, when there is no multiplexing, this measure would be $1/L$.
@@
High Multiplexing X Seed Set Centrality $-0.039$
@@ 3.1. A Model of Diffusion with Multiplexing
We study diffusion/contagion in a society consisting of a finite set of individuals $N=\{1,\ldots,n\}$. Each individual has relationships captured via layers $\{1,\ldots,L\}$, with a generic layer represented by $\ell$. In each layer $\ell$, the interactions between individuals are described by a (possibly directed) network with adjacency matrix $g^{\ell}\in\{0,1\}^{n\times n}$, such that $g^{\ell}_{ij}=1$ if there is a link from $i$ to $j$ in layer $\ell$ (interpreted as being capable of being infected by $j$, e.g., via $i$ paying attention to $j$ in a model of information flow), and 00 otherwise. We denote the multigraph consisting of $L$ layers by $g=(g^{1},g^{2},\ldots,g^{L})$.
@@
At each time $t$, an individual’s state can change based on the infection status of its neighbors. A susceptible individual $i$ becomes infected if it receives at least $\tau$ infection transmissions from its infected neighbors in a given time period. An infected individual recovers (and becomes susceptible again) randomly with a probability $\delta$ at the end of a period. If $\tau=1$, this represents a standard (simple) contagion process, while with a threshold $\tau>1$ this is known as a complex contagion (Granovetter, 1978; Centola, 2010). [^14] This is closely related to games on networks (Morris, 2000; Jackson and Zenou, 2014).
@@
This is the probability of $k$ transmissions; note $f(k;\mathcal{L}_{ij})$ can capture arbitrary patterns of correlation in infection transmission through multiple layers. For each layer $\ell$, let $q_{\ell}\in(0,1)$ be the marginal probability of infection transmission from an infected individual to a susceptible one if they are connected via that layer. We allow different layers to have different contact probabilities, which is needed given the heterogeneity in the roles of different layers discussed in Section 2.3. If there is a positive correlation in transmission across layers, two nodes connected by layers $\{A,B\}$ have an infection distribution satisfying $f(2,\{A,B\})\geq q_{A}q_{B}$.
The probability that a susceptible individual $i$ becomes infected at time $t$ given the infection status of its neighbors at time $t-1$ is
@@ 3.1.1. Comparisons of Multiplexing
To formalize this type of ranking, we define a local multiplexity dominance relation, denoted by $\prec$. For two multigraphs $g$ and $\widehat{g}$, we say $\widehat{g}\prec g$—that is $\widehat{g}$ is locally less multiplexed than $g$—if $\widehat{g}$ can be obtained from $g$ by removing a link in some layer $\ell$ between nodes $i$ and $j$ and adding a new link in that same layer to another neighbor $k$, where $i$’s connections to $k$ occurred in a set of layers that form a strict subset of the layers (except layer $\ell$) in which $i$ was connected with $j$ to start with. This means that: (i) $\mathcal{L}_{ik}(g)\subsetneq\mathcal{L}_{ij}(g)\setminus\{\ell\}$, (ii) ${g}_{ik}^{\ell}=0=\widehat{g}_{ij}^{\ell}$ and $\widehat{g}_{ik}^{\ell}=1={g}_{ij}^{\ell}$, and (iii) for all other links $\widehat{g}$ and $g$ coincide.
@@ 3.2.1. Infection of an Individual
where $A,B$ are the layers. Simplifying this yields
@@
Consider simple contagion ($\tau=1$). If $\widehat{g}\overline{\prec}_{i}g$ and each of $i$’s neighbors is infected independently with probability $\rho>0$, and $i$ is susceptible, then $i$ is more likely to be infected under the less multiplexed network $g$ than under $\widehat{g}$ if and only if transmission is not too negatively correlated across layers (condition 3.1), with the reverse holding if condition 3.1 fails.
@@ 3.2.2. Multiplexing and Overall Infection in the SIS Model
Focusing again on the case of two layers, a sufficient statistic for $D_{i}$ for the mean-field analysis is a triple $\hat{D}_{i}=(\hat{D}_{iA},\hat{D}_{iB},\hat{D}_{i,AB})$, which represents the number of connections that $i$ has that are just on layer $A$, just on layer $B$, and on both layers, respectively. The distribution of $\hat{D}$ across the population is described by a function $P(\hat{D})$ that has finite support. The steady-state infection rate of nodes with connection profile $\hat{D}$ is denoted $\rho(\hat{D})$. The population infection rate is then defined by
(3.2) $\rho=\sum_{\hat{D}}P(\hat{D})\rho(\hat{D}).$
The probability that a susceptible node with connections $\hat{D}=(\hat{D}_{A},\hat{D}_{B},\hat{D}_{AB})\neq(0,0,0)$ becomes infected, in steady state, is then [^17] Here, $(1-\rho q_{A})$ is the probability that an infected layer-$A$-only neighbor fails to transmit infection, and similarly for $(1-\rho q_{B})$ for a layer-$B$-only neighbor. For neighbors connected via both layers, $(1-\rho[q_{A}+q_{B}-f(2,\{A,B\})])$ is the probability that such a neighbor fails to transmit infection, accounting for potential correlation in transmissions across the two layers. Raising these terms to the powers $\hat{D}_{A},\hat{D}_{B},\hat{D}_{AB}$ accounts for all relevant neighbors. Multiplying them together gives the probability that none of these neighbors transmit infection through their respective sets of layers. Subtracting this product from 1 then yields the probability that at least one transmission succeeds, infecting the susceptible node.
@@
A steady-state is a joint solution to (3.2) and (3.4) for each $\hat{D}$ in the support of $P$. Note that 00 is always a solution, and for some distributions $P$ there may also exist a positive solution. We focus on the largest positive solution, which is the one that corresponds to the behavior of large finite graphs. [^18] See Elliott et al. (2022) for a detailed argument in an analogous situation.
We extend the partial order we defined in 3.1.1 to the space of distributions as follows. We say that a distribution ${P}^{\prime}$ is less multiplexed that $P$, denoted by $P^{\prime}\overline{\prec}P$, if there exists $\hat{D}$ and $\hat{D}^{\prime}$ such that
@@
Consider a simple contagion process ($\tau=1$) process. Let transmission probabilities be given by $f$ with marginal probabilities $(q^{\ell})_{\ell}\in(0,1)^{L}$. Finally, fix a recovery rate $\delta\in(0,1)$ and two distributions of connections $P^{\prime}$ and $P$ that each have positive steady-state infection rates. If $P^{\prime}\overline{\prec}P$, then the positive steady-state infection rate under $P^{\prime}$ is higher than that under $P$ if and only if transmission is not too negatively correlated (condition 3.1) at the positive infection rate of $P$.

### v2

@@ 2.3. Correlations across Network Layers
First, there are consistently high correlations between layers in both data sets—above 0.5 for most layer pairs. Second, the exceptions are the distance, jati, and temple layers. The jati and distance layers are almost uncorrelated with the other layers, [^4] This does not mean, for instance, that there is not substantial jati-based homophily in these data. The low correlation comes from the fact that the jati layer dramatically over-predicts relationships compared to other layers, so it has many 1’s where there are 0’s in the other layers., [^5] Distance is higher when people live far from each other and are thus less likely to be linked, all else held equal; this explains the negative signs. while the temple layer has an intermediate level of correlation with others. Third, the layers are more highly correlated in the RCT villages compared to the microfinance villages.
2.4. Principal Components and the Backbone Network
To further analyze the network structure across layers, we apply a principal component analysis (PCA) to the multigraph of each village. We perform the analysis with all of the layers (excluding the synthetic union and intersection layers). We treat each pair of households (in a given village) as an observation, yielding $\sum_{v}\binom{n_{v}}{2}$ observations, where $n_{v}$ is the number of households in village $v$, and the number of dimensions is the number of layers $L$ in the given sample. (Details for how the PCA was implemented can be found in Appendix Section A.3.)
@@ 2.4. Principal Components and the Backbone Network
The first principal component aligns closely with most relationship layers and captures 4949% of variation in the microfinance sample and 7272% in the RCT sample (Fig 2 panels A and B). Jati, by contrast, loads primarily on the second component, consistent with its limited correlation with other ties. Geographic distance is negatively aligned with jati, reflecting residential clustering by caste. Complete loadings for each layer appear in Supplementary Tables S3 and S4.
Next, in Figure 3 we repeat the analysis after removing the least correlated dimensions: jati, geography, and temple. [^6] We also redo the analysis just dropping jati and geography and keeping temple in Supplemental Appendix Figure S2. Temple is sparse and essentially orthogonal to the other dimensions. This allows us to zoom in on the correlation patterns among the social and economic layers.
@@
To formally capture the correlation structure of these network layers, we use the principal components to construct a continuous weighted network, which we call the backbone. The backbone network is built using the first $K$ principal components, where the optimal $K$ is derived from the so-called ladle plot from Luo and Li (2016) (see Appendix Section A.3 for details). The backbone provides a low-dimensional index which reduces the multiplex data to a synthetic structure by projecting multidimensional links onto the top principal components.
3. Multiplexing and Information Diffusion
@@
We use a specific diffusion centrality measure developed in Banerjee et al. (2013) and further studied in Banerjee et al. (2019). The details of that measure appear in Appendix Section A.4.
@@
Table 1 depicts how differently the layers predict diffusion based on our specification in (3.1). (Supplementary Appendix Figure S1 plots the 9090% and 9595% confidence intervals.)
@@
Interestingly, the four synthetic networks we have mentioned that aggregate the layers in specific ways—union, intersection, total, and backbone—all perform worse than the individual layers with the exception of jati. However, this appears to be rooted in the inclusion of jati in those aggregates. In Supplement C.1 we recreate Table 1 with aggregate layers that omit jati in their construction (this applies only to union, intersection, and backbone). This improves their performance, with the backbone network now yielding an $R^{2}$ second only to the advice layer.
@@
(3.2) $y_{v}=\alpha+\sum_{\ell}\beta^{\ell}\cdot DC_{v}^{\ell}+X_{v}\Gamma+\epsilon_{v}$
@@
A complication we face here is that in order to be consistent, LASSO requires a condition called irrepresentability, which requires the regressors of interest not to be excessively correlated (Zhao and Yu, 2006). In our setting, this requirement fails since the network layers are highly correlated. To overcome this problem, we use the Puffer transformation developed by Rohe (2015) and Jia et al. (2015), which recovers irrepresentability when the number of observations exceeds the number of variables. See Appendix section A.5 for details.
@@
  • Note: We present both cumulative and marginal F-tests as layers are successively added in the order selected by LASSO. The “F-stat” and “p-val” columns correspond to the cumulative test comparing each specification with the intercept only benchmark. “F-stat marginal” and “p-val marginal” columns correspond to the marginal test when adding a given layer.
Table 3 presents both cumulative and marginal F-tests as variables are added in the order selected by LASSO. We can see that adding intersection is marginally significant above advice, and further including both kerorice and jati together yields a more complete model, with an improvement significant at the 5 percent level. [^7] In Appendix Table S5 we exclude the extra layers of intersection, union, total links and backbone, which are “constructed” layers that are derived from these basic layers. F-tests include the basic layers in the order selected by the Lasso. Thus, even though jati serves as a poor substitute for other layers, it turns out to be a useful complement to them in predicting diffusion.
@@ 3.1. How the Level of Multiplexing Affects Diffusion
The multiplexing score for a household $i$ measures the average fraction of relationship types it has with each of its neighbors. The numerator is a sum over all neighbors $j$ of the fraction of layers on which household $i$ is connected to $j$. Specifically, for each neighbor $j$, it calculates the fraction of layers connecting $i$ to $j$ by summing the number of links between them across all layers and dividing by the total number of layers $L$, then sums these fractions across all neighbors. The denominator counts the number of unique neighbors of household $i$ by summing an indicator for whether there is at least one link between $i$ and $j$ across any layer. The overall score is thus the average, across neighbors, of the fraction of layers shared with each neighbor. For example, $m_{i,v}=1$ if whenever household $i$ has a relationship with some other household $j$, then it has all possible relationships with that other household. In contrast, when there is no multiplexing, this measure would be $1/L$.
@@
High Multiplexing X Seed Set Centrality 0.039-0.039
@@ 4.1. A Model of Diffusion with Multiplexing
We study diffusion/contagion in a society consisting of a finite set of individuals $N=\{1,\ldots,n\}$. Each individual has relationships captured via layers $\{1,\ldots,L\}$, with a generic layer represented by $\ell$. In each layer $\ell$, the interactions between individuals are described by a (possibly directed) network with adjacency matrix $g^{\ell}\in\{0,1\}^{n\times n}$, such that $g^{\ell}_{ij}=1$ if $j$ can transmit to $i$, and 0 otherwise. (We still speak of the associated edge in the directed graph as being directed from $i$ to $j$, which can be thought of, e.g., as $i$ paying attention to $j$.) We denote the multigraph consisting of $L$ layers by $g=(g^{1},g^{2},\ldots,g^{L})$.
@@
At each time $t$, an individual’s state can change based on the infection status of its neighbors. A susceptible individual $i$ becomes infected if it receives at least $\tau$ infection transmissions from its infected neighbors in a given time period. An infected individual recovers (and becomes susceptible again) randomly with a probability $\delta$ at the end of a period. If $\tau=1$, this represents a standard (simple) contagion process, while with a threshold $\tau>1$ this is known as a complex contagion (Granovetter, 1978; Centola, 2010). [^8] This is closely related to games on networks (Morris, 2000; Jackson and Zenou, 2014). The greater threshold captures that some behaviors may require more thought, reinforcement, or interactions to become attractive. For background evidence for such thresholds, there are many such applications as detailed in Centola (2010, 2011, 2018); Boucher et al. (2024), for example.
@@
This is the probability of $k$ transmissions; note $f(k;\mathcal{L}_{ij})$ can capture arbitrary patterns of correlation in infection transmission through multiple layers. For each layer $\ell$, let $q_{\ell}\in(0,1)$ be the marginal probability of infection transmission from an infected individual to a susceptible one if they are connected via that layer.
We emphasize that we allow different layers to have different contact probabilities, which is needed given the heterogeneity in the roles of different layers discussed in Section 3. One can hear information about a new loan program when stopping by to borrow rice from a friend, but it could be more likely that the friend will mention it if one is stopping by to ask for advice. It could also be that there are two conversations about the loan program, and these could be correlated. If there is a positive correlation in transmission across layers, two nodes connected by layers $\{A,B\}$ have an infection distribution satisfying $f(2,\{A,B\})\geq q_{A}q_{B}$, with negative correlation being the opposite.
The probability that a susceptible individual $i$ becomes infected at time $t$ given the infection status of its neighbors at time t1t-1 is
@@ 4.1.1. Comparisons of Multiplexing
To formalize this type of ranking, we define a local multiplexity dominance relation, denoted by $\prec$. For two multigraphs $g$ and $\widehat{g}$, we say $\widehat{g}\prec g$—that is $\widehat{g}$ is locally less multiplexed than gg—if $\widehat{g}$ can be obtained from $g$ by removing a link in some layer $\ell$ between nodes $i$ and $j$ and adding a new link in that same layer to another neighbor $k$, where $i$’s connections to $k$ occurred in a set of layers that form a strict subset of the layers (except layer $\ell$) in which $i$ was connected with $j$ to start with. This means that: (i) $\mathcal{L}_{ik}(g)\subsetneq\mathcal{L}_{ij}(g)\setminus\{\ell\}$, (ii) ${g}_{ik}^{\ell}=0=\widehat{g}_{ij}^{\ell}$ and $\widehat{g}_{ik}^{\ell}=1={g}_{ij}^{\ell}$, and (iii) for all other links $\widehat{g}$ and $g$ coincide.
@@ 4.2.1. Infection of an Individual
where $A,B$ are the layers. [^10] On the left side, $q_{A}\rho+q_{B}\rho$ represents infection from the single neighbor $j$ via either layer, and we subtract the overlap where both layers transmit. On the right, the same sum represents infection from two distinct neighbors ($j$ via layer $A$ or $k$ via layer $B$), and we subtract the probability that both transmit independently. Simplifying this yields
@@
Consider simple contagion ($\tau=1$). If $\widehat{g}\;\overline{\prec}_{i}\;g$ and each of $i$’s neighbors is infected independently with probability $\rho>0$, and $i$ is susceptible, then $i$ is more likely to be infected under the less multiplexed network $\widehat{g}$ than under $g$ if and only if transmission is not too negatively correlated across layers (condition 4.1), with the reverse holding if condition 4.1 fails.
@@ 4.2.2. Multiplexing and Overall Infection in the SIS Model
[…]
The probability that a susceptible node with connections $\hat{D}=(\hat{D}_{A},\hat{D}_{B},\hat{D}_{AB})\neq(0,0,0)$ becomes infected, in steady state, is then [^12] Here, $(1-\rho q_{A})$ is the probability that a layer-$A$-only neighbor fails to transmit infection, and similarly for $(1-\rho q_{B})$ for a layer-$B$-only neighbor. For neighbors connected via both layers, $(1-\rho[q_{A}+q_{B}-f(2,\{A,B\})])$ is the probability that such a neighbor fails to transmit infection, accounting for potential correlation in transmissions across the two layers. Raising these terms to the powers $\hat{D}_{A},\hat{D}_{B},\hat{D}_{AB}$ accounts for all relevant neighbors. Multiplying them together gives the probability that none of these neighbors transmit infection through their respective sets of layers. Subtracting this product from 1 then yields the probability that at least one transmission succeeds, infecting the susceptible node.
@@
A steady-state is a joint solution to (4.2) and (4.4) for each $\hat{D}$ in the support of $P$. Note that 0 is always a solution, and for some distributions $P$ there may also exist a positive solution. We focus on the largest positive solution, which is the one that corresponds to the behavior of large finite graphs. [^13] See Elliott et al. (2022) for a detailed argument in an analogous situation.
We extend the partial order we defined in 4.1.1 to the space of distributions as follows. We say that a distribution ${P}^{\prime}$ is less multiplexed than $P$, denoted by $P^{\prime}\overline{\prec}P$, if there exists $\hat{D}$ and $\hat{D}^{\prime}$ such that
@@
Consider a simple contagion process ($\tau=1$) process. Let transmission probabilities be given by $f$ with marginal probabilities $(q^{\ell})_{\ell}\in(0,1)^{L}$. Finally, fix a recovery rate $\delta\in(0,1)$ and two distributions of connections $P^{\prime}$ and $P$ that each have positive steady-state infection rates. If $P^{\prime}\overline{\prec}P$, then the positive steady-state infection rate under $P^{\prime}$ is higher than that under $P$ if and only if transmission is not too negatively correlated (condition 4.1) at the positive infection rate of $P$.